OpenStudy (anonymous):

Find all solutions to the system using the Gauss-Jordan elimination algorithm. (system soon to come)

OpenStudy (anonymous):

\[x_1+2x_2+2x_3=6\]\[2x_1+x_2+x_3=6\]\[x_1+x_2+3x_3=6\]